OpenStudy (hippoharry):

The vertex form of the equation of a parabola is y = (x + 7)2 + 13. What is the standard form of the equation? A. y = 7x2 + 14x+ 62 B. y = x2 + 7x + 13 C. y = x2 + 43x + 32 D. y = x2 + 14x + 62

OpenStudy (mathmale):

Please, use " ^ " to indicate exponentiation. If you are "cutting and pasting" problems, you lose the " ^ ". You must put it in manually.

OpenStudy (mathmale):

y = (x + 7)^2 + 13 is the same as \[y = (x + 7)^2 + 13\]

OpenStudy (hippoharry):

The vertex form of the equation of a parabola is y = (x + 7)^2 + 13. What is the standard form of the equation? A. y = 7x^2 + 14x+ 62 B. y = x^2 + 7x + 13 C. y = x^2 + 43x + 32 D. y = x^2 + 14x + 62

OpenStudy (mathmale):

Compare this to y=a(x-h)^2+k. What are your h and k? The point (h,k) is the vertex of your parabola.